# Characterization of NBI-driven shear Alfvén waves in the TJ-II stellarator using Mirnov probes and electrostatic potential fluctuation measurements

### v1

@@ 5 Time evolution of rotational transform
To understand the changes that occur in the Alfvén wave spectrum when we modify the direction of the microwaves beam it is necessary to have an estimate of the time evolution of the rotational transform. When this evolution is due to a current source that is established from almost the beginning or when we observe that the value of the toroidal current measured by the Rogowski coil stabilizes before the end of the shot we can assume that, in the final part of the shot, the plasma shielding current is zero and the current profile inside the plasma is that of the current source that we force in the plasma, either using ECCD or NBCD, or is the one produced by the bootstrap current inherent to the configuration and the plasma profiles. With this approximation we can have an estimate of the rotational transform at the end of the shot [cappaStabilityAnalysis2021]. In this case this cannot be done because we have a source of ECCD current during the whole shot duration, and a second source due to the NBI which is established a few milliseconds after the neutral beam injection. Furthermore we observe in most of the cases that the current is not stabilized at the end of the shot.
@@ 5.1 Radial diffusion of toroidal current
Equation 6 is equivalent to the one given for example in [zaniniECCDinducedSawtooth2020] taking into account that the relation between the current and iota in the cylindrical case is given by
@@ 5.2 Initial and boundary conditions
As for the initial conditions, we stated above that the main current sources come from ECCD and NBCD, that is . For both the ECCD current at different injection angles and the NBCD current produced by the counter injector we will use a combination of estimates presented in previous works combined with the plasma toroidal current measured in the experiments discussed here.
@@
The NBCD profile has been taken from the ASCOT5 slowing-down simulations presented in [mulasASCOT5Simulations2022]. The total NBCD current obtained in the medium configuration was validated against experimental results in [mulasValidatingNeutralbeam2023] and a good agreement was observed for neutral beam counter-injection at line densities and density profiles similar to the ones shown in figures 4 and 6. For our present purposes and having in mind that different configurations are used and that the plasma densities slightly differ between the three of them, the integrated current of the NBCD source will be determined from the plasma toroidal current measured in the cases and we will take the current density profile calculated in reference [mulasValidatingNeutralbeam2023]. We mentioned in section 2 that the low frequency oscillations originated by inductive coupling with the main conductors can be extracted from the Rogowski coil measurement. Note that these oscillations not only occur with respect to a certain average value, but may also be shifted upwards or downwards depending on the specific behavior of current ripple in the central and helical coils of the device. Applying the procedure described in the appendix of [mulasValidatingNeutralbeam2023] to the current measured in the NBI phase we arrive at the result shown in figure 22.
@@
The values of the integrated NBCD current source (no ECCD in this case) in the limit of zero shielding current and the LR time for each case are given in the figure. The current measured in the medium and high iota configurations (actually the limit current inferred from the measurements) agrees well with the results presented in [mulasValidatingNeutralbeam2023]. However, the NBCD measured in the low iota configuration appears to be lower. A possible reason for this is the slightly lower plasma volume and size of this configuration (see figure 3) and also the fact that plasma line densities in this case () were a bit lower than the line densities achieved in the other two (). The beam shine-trough power and subsequent slowing-down ASCOT5 simulations that would clarify these differences are out of the scope of this work. The ECCD profile is estimated from experimental observations of ECCD total induced current in oblique launching experiments [fernandezECCDExperiments2008] and fast power modulation experiments [eguiliorHeatWave2003] for perpendicular injection combined with the results of ray-tracing and beam tracing simulations [martinez-fernandezHighPower2020]. We use these simulations to determine how the power deposition and ECCD profiles widens as increases. This result is corrected with the measured power deposition profile at by imposing that its integral is the total measured current after several LR times. With these assumptions we obtain the current densities that are shown in figure 23.
@@ 5.3 Iota evolution results without ECCD
The calculated evolution shows that for these cases in which the toroidal current is still evolving (see figure 22), the profile of at the end of the shot is midway between the vacuum value and the equilibrium value. It is enlightening to compare the profiles shown in the figures with the mode numbers of the low frequency MHD instabilites ( kHz) shown in figure 15. Restricting ourselves to the shots with for the moment, the mode number analysis shows a instability for the low configuration, a for the medium case and two instabilites with mode numbers and for the high case. This is consistent with the calculated profiles considering that this type of instability is triggered in the density gradient region () in the presence of a low order rational value; , and , for the low, medium and high iotabar respectively. Note that following the numerical study of Mirnov coils performance presented in [pons-villalongaExploringOperational2024] the calculation of the mode numbers is subject to an indeterminacy of or greater depending on the radial position of the instability and thus the pairs and delivered by the analysis may very well correspond to and having in mind that actually and are well outside the reasonable values when negative NBCD current is at play.
@@ 5.4 Iota evolution results with ECCD
For maximum positive ECCD (), disappears since positive current elevates the profile. The most interesting case, for which the calculation of the time evolution of iota may give us a clue, is the one for . In this case, co-ECCD pushes-up the profile but at soon as NBCD comes into play, its contribution makes the rational surface appear again. This is also the case for which the perturbation shows the largest amplitude in the magnetic signal. The two-step model cannot capture exactly the time sequence of the experiment but clearly illustrates the phenomenon. This is shown in 25, where both the final iota profile due to ECCD current only and the final profile due to the combination of both sources is presented. The first one does not contain the low order rational while the second one intersects reaches at and .
@@ 5.5 MSE measurements
Using the VMEC code to calculate the plasma equilibrium and its profile for a given profile of toroidal current, and determining the components of the field perpendicular to the beam, we may estimate the quantity measured by MSE and establish a correspondance with the profile. Figure 28 shows the results of the MSE measurements for different launching angles of the ECRH1 beam. These were made at the end of the NBI+ECRH phase where both NBCD and ECCD are acting to modify the rotational transform. Although MSE can clearly distinguish between rotational transforms of different configurations [mccarthySpectrallyResolved2015], it cannot resolve changes in due only to different plasma currents.
@@ 6 Shear Alfvén continuum with STELLGAP
Figure 31 shows the result of the STELLGAP calculation for . Each plot corresponds to one mode family. The sign criterion for the phase term in STELLGAP () is opposite to the one used in the Lomb periodogram analysis and therefore must be replaced by when comparing the results of the code to the measurements. Two helical gaps (HAE21 and HAE31) in the spectrum appear for mode families and . TAE gap appears higher in frequency and is not shown here. According to the result shown in 15, the 9/4 mode, observed around 175 GHz, would correspond to the {, } pair in the HAE31 gap. Previous studies with a synthetic Mirnov diagnostic show [pons-villalongaExploringOperational2024] that, besides uncertainties in poloidal mode number determination up to , Mirnov coils arrays in TJ-II cannot distinguish different constituents in helically coupled structures and therefore we cannot rule out that such combination of modes may show as an when measured. Other modes observed in the case, identified with less degree of confidence as and could have its correspondance in the family modes. In this case, the minimum in frequency of the mode favors the presence of a global Alfven eigenmode (GAE) gap while the mode would have to be an energetic particle mode (EPM) since it would intersect the continuum. None of this can be assured with certainty without calculating the profiles of the modes, using a code such as AE3D, and their destabilization rates by means of linear stability simulations using gyro-fluid or gyrokinetic codes such as FAR3d or EUTERPE respectively. In the same way we can take now the rotational transform profile of the case and calculate the corresponding continuum.
The minimum in the rotational transform profile around , produced by the combination of positive ECCD and negative NBCD has a strong impact of the shear Alfvén spectrum as we could already guess from the experimental observation that shows up to five different modes in the magnetic fluctuations spectrum (see figure 14). Modes with low mode numbers , , determined with a high degree of reproducibility in the periodogram analysis, and very close in frequency are consistemt with the continuum of the family shown in figure 32. There, maximums and minimums in the continuum, also close in frequency, and prone to GAE destabilization appear around with mode numbers and . Moreover, a high frequency mode is observed around 310 kHz, where the calculation predicts also high mode numbers AEs in the HAE21 and HAE31 gaps that merge in the plasma periphery. HIBP measurements for these modes are shown in figure 19.
@@ 7 Conclusions
We have described a series of experiments performed in TJ-II in which ECCD was used to modify the profile in order to destabilize different Alfvén eigenmodes and to have a wide range of cases available for further validation studies. The observed modes have been diagnosed with two arrays of Mirnov coils, used to determine the toroidal and poloidal mode numbers, and one heavy ion beam probe operated in radial sweep mode to measure the radial profile of the perturbation. This is the first time that the toroidal mode number of Alvén waves in TJ-II has been measured. Significantly, radial asymmetries were observed in the potential profiles of some modes, which points to their character as gap modes. There remain uncertainties in the mode identification, some due to the inherent limitations of the diagnostic and others that could be mitigated with an improved analysis method.
For the purpose of theory validation, the determination of the radial profile of the rotational transform is essential. It is shown that changes in iota that can lead to significant variations in the Alfvén continuum are not measurable with MSE and, instead, a cylindrical model for the evolution of the shielding current in response to the NBI- and EC-induced currents has been used to estimate it. The profiles obtained with this model have been used as inputs for STELLGAP, and the results show a reasonable consistency between the calculated shear Alfvén continua and the experimental observations.
The database of AEs obtained in different scenarios, characterized by their mode numbers and radial profiles of electric potential can be used for theory validation purposes using gyro-fluid codes such as FAR3d [varelaStabilityOptimization2024], already applied to co-NBI driven modes in TJ-II [cappaStabilityAnalysis2021] or gyrokinetic codes such as EUTERPE [kleiberEUTERPEGlobal2024], that was applied to study AEs fast-ion drive in W7-X [slabyInvestigationMode2020]. Thermal plasma profiles and estimates of neutral density profiles are available and fast ion slowing down simulations with ASCOT5 in the three magnetic configurations can be performed to obtain the corresponding fast ion density profiles. In addition to mode potential, HIBP measurements can also provide the density fluctuations profile associated to the Alfvén mode [melnikov2DDistributions2022], which are very relevant when comparing with the theoretical predictions.

### v2

@@ 5 Time evolution of rotational transform
To understand the changes that occur in the Alfvén wave spectrum when we modify the direction of the microwaves beam it is necessary to have an estimate of the time evolution of the rotational transform. When this evolution is due to a current source that is established from almost the beginning or when we observe that the value of the toroidal current measured by the Rogowski coil stabilizes before the end of the shot we can assume that, in the final part of the shot, the plasma shielding current is zero and the current profile inside the plasma is that of the current source that we force in the plasma, either using ECCD or NBCD, or is the one produced by the bootstrap current inherent to the configuration and the plasma profiles. With this approximation we can have an estimate of the rotational transform at the end of the shot [cappaStabilityAnalysis2021]. In this case this cannot be done because we have a source of ECCD current during the whole shot duration, and a second source due to NBI2 which is established a few milliseconds after the neutral beam injection. Furthermore we observe in most of the cases that the current is not stabilized at the end of the shot.
@@ 5.1 Radial diffusion of toroidal current
Equation 6 is equivalent to the one given in [zaniniECCDinducedSawtooth2020] taking into account that the relation between the current and iota in the cylindrical case is given by
@@ 5.2 Initial and boundary conditions
As for the initial conditions, we stated above that the main current sources come from ECCD and NBCD, that is . For both the ECCD current at different injection angles and the NBCD current produced by the counter injector we will use a combination of estimates presented in previous works together with the plasma toroidal current measured in the experiments discussed here.
@@
The NBCD profile has been taken from the ASCOT5 slowing-down simulations presented in [mulasASCOT5Simulations2022]. The total NBCD current obtained in the medium configuration was validated against experimental results in [mulasValidatingNeutralbeam2023] and a good agreement was observed for neutral beam counter-injection at line densities and density profiles similar to the ones shown in figures 4 and 6. For our present purposes and having in mind that different configurations are used and that the plasma densities slightly differ between the three of them, the integrated current of the NBCD source will be determined from the plasma toroidal current measured in the cases and we will take the current density profile calculated in reference [mulasValidatingNeutralbeam2023]. We mentioned in section 2 that the low frequency oscillations originated by inductive coupling with the main conductors can be extracted from the Rogowski coil measurement. Note that these oscillations not only occur with respect to a certain average value, but may also be shifted upwards or downwards depending on the specific behavior of current ripple in the central and helical coils of the device. Applying the procedure described in the appendix of [mulasValidatingNeutralbeam2023] to the current measured in the NBI2 phase we arrive at the result shown in figure 22.
@@
The values of the integrated NBCD current source (no ECCD in this case) in the limit of zero shielding current and the LR time for each case are given in the figure. The current measured in the medium and high iota configurations (actually the limit current inferred from the measurements) agrees well with the results presented in [mulasValidatingNeutralbeam2023]. However, the NBCD measured in the low iota configuration appears to be lower. A possible reason for this is the slightly lower plasma volume and size of this configuration (see figure 3) and also the fact that plasma line densities in this case () were a bit lower than the line densities achieved in the other two (). The beam shine-trough power and subsequent slowing-down ASCOT5 simulations that would clarify these differences are out of the scope of this work.
The ECCD profile is estimated from experimental observations of ECCD total induced current in oblique launching experiments [fernandezECCDExperiments2008] and fast power modulation experiments [eguiliorHeatWave2003] for perpendicular injection combined with the results of ray-tracing and beam tracing simulations [martinez-fernandezHighPower2020]. We use these simulations to determine how the power deposition and ECCD profiles widens as increases. This result is corrected with the measured power deposition profile at by imposing that its integral is the total measured current after several LR times. With these assumptions we obtain the current densities that are shown in figure 23.
@@ 5.3 Iota evolution results without ECCD
The calculated evolution shows that for these cases in which the toroidal current is still evolving (see figure 22), the profile of at the end of the shot ( ms) is midway between the vacuum value and the equilibrium value. It is enlightening to compare the profiles shown in the figures with the mode numbers of the low frequency MHD instabilites ( kHz) shown in figure 15. Restricting ourselves to the shots with for the moment, the mode number analysis shows a instability for the low configuration, a for the medium case and two instabilites with mode numbers and for the high case. This is consistent with the calculated profiles considering that this type of instability is triggered in the density gradient region () in the presence of a low order rational value; , and , for the low, medium and high respectively. Note that following the numerical study of Mirnov coils performance presented in [pons-villalongaExploringOperational2024] the calculation of the mode numbers is subject to an indeterminacy of or greater depending on the radial position of the instability and thus the pairs and delivered by the analysis may very well correspond to and having in mind that actually and are well outside the reasonable values when negative NBCD current is at play.
@@ 5.4 Iota evolution results with ECCD
For maximum positive ECCD (), disappears since positive current elevates the profile. The most interesting case, for which the calculation of the time evolution of iota may give us a clue, is the one for . In this case, co-ECCD pushes-up the profile but at soon as NBCD comes into play, its contribution makes the rational surface appear again. This is also the case for which the perturbation shows the largest amplitude in the magnetic signal. The two-step model cannot capture exactly the time sequence of the experiment but clearly illustrates the phenomenon. This is shown in 25, where both the final iota profile due to ECCD current only and the final profile due to the combination of both sources is presented. The first one does not contain the low order rational while the second one intersects at and .
@@ 5.5 MSE measurements
Using the VMEC code to calculate the plasma equilibrium and its profile for a given profile of toroidal current, and determining the components of the field perpendicular to the beam, we may estimate the quantity measured by MSE and establish a correspondance with the profile. Figure 28 shows the results of the MSE measurements for different launching angles of the ECRH1 beam. These were made at the end of the NBI2+ECRH(1&2) phase where both NBCD and ECCD are acting to modify the rotational transform. Although MSE can clearly distinguish between rotational transforms of different configurations [mccarthySpectrallyResolved2015], it cannot resolve changes in due only to different plasma currents.
@@ 6 Shear Alfvén continuum with STELLGAP
Figure 31 shows the result of the STELLGAP calculation for . Each plot corresponds to one mode family. The sign criterion for the phase term in STELLGAP () is opposite to the one used in the Lomb periodogram analysis and therefore must be replaced by when comparing the results of the code to the measurements. Two helical gaps (HAE21 and HAE31) in the spectrum appear for mode families and . TAE gap appears higher in frequency and is not shown here. According to the result shown in figure 15, the 9/4 mode, observed around 175 GHz, would correspond to the {, } pair in the HAE31 gap. Previous studies with a synthetic Mirnov diagnostic [pons-villalongaExploringOperational2024] show that, besides uncertainties in poloidal mode number determination up to , Mirnov coils arrays in TJ-II cannot distinguish different constituents in helically coupled structures and therefore we cannot rule out that such combination of modes may show as an when measured. Other modes observed in the case, identified with less degree of confidence as and could have its correspondance in the family modes. In this case, the minimum in frequency of the mode favors the presence of a global Alfvén eigenmode (GAE) while the mode would have to be an energetic particle mode (EPM) since it would intersect the continuum. None of this can be assured with certainty without calculating the profiles of the modes, using a code such as AE3D, and their destabilization rates by means of linear stability simulations using gyro-fluid or gyrokinetic codes such as FAR3d or EUTERPE respectively.
In the same way we can take now the rotational transform profile of the case and calculate the corresponding continuum.
The minimum in the rotational transform profile around , produced by the combination of positive ECCD and negative NBCD has a strong impact of the shear Alfvén spectrum as we could already guess from the experimental observation that shows up to five different modes in the magnetic fluctuations spectrum (see figure 14). Modes with low mode numbers , , determined with a high degree of reproducibility in the periodogram analysis, and very close in frequency are consistent with the continuum of the family shown in figure 32. There, maximums and minimums in the continuum, also close in frequency, and prone to GAE destabilization appear around with mode numbers and . Moreover, a high frequency mode is observed around 310 kHz, where the calculation for predicts also high mode numbers AEs in the HAE21 and HAE31 gaps that merge in the plasma periphery. HIBP measurements for these modes are shown in figure 19.
@@ 7 Conclusions
We have described a series of experiments performed in the TJ-II stellarator. ECCD was used to modify the profile in order to destabilize different NBI-driven Alfvén eigenmodes and to have a wide range of cases available for further validation studies. The observed modes have been diagnosed with two arrays of Mirnov coils, used to determine the toroidal and poloidal mode numbers, and one heavy ion beam probe operated in radial sweep mode to measure the radial profile of the perturbation. This is the first time that the toroidal mode number of Alvén waves in TJ-II has been measured. Significantly, radial asymmetries were observed in the potential profiles of some modes, which points to their character as gap modes. There remain uncertainties in the mode number measurements, some due to the inherent limitations of the diagnostic and others that could be mitigated with an improved analysis technique, such as the SSI method [kleiberModernMethods2021], or Dynamic Mode Decomposition (DMD) [schmidDynamicMode2010].
For the purpose of theory validation, the determination of the radial profile of the rotational transform is essential. It is shown that changes in iota that can lead to significant variations in the Alfvén continuum are not measurable with MSE and, instead, a cylindrical model for the evolution of the shielding current in response to the NBI- and EC-induced currents has been used to estimate it. The profiles obtained with this model have been used as inputs for STELLGAP, and the results show a reasonable consistency between the calculated shear Alfvén continua and the experimental observations. Only the cases with the most intense activity have been analyzed with STELLGAP, and other interesting cases, such as the one exhibiting a structure with three very close frequencies and a clearly asymmetric radial profile, all compatible with gap mode behavior, have been left for a more theoretical treatment.
The database of AEs obtained in different scenarios, characterized by their mode numbers and radial profiles of electrostatic potential can be used for theory validation purposes using gyro-fluid codes such as FAR3d [varelaStabilityOptimization2024], already applied to co-NBI driven modes in TJ-II [cappaStabilityAnalysis2021] or gyrokinetic codes such as EUTERPE [kleiberEUTERPEGlobal2024], that was applied to study AEs fast-ion drive in W7-X [slabyInvestigationMode2020]. Thermal plasma profiles and estimates of neutral density profiles are available and fast ion slowing down simulations with ASCOT5 in the three magnetic configurations can be performed to obtain the corresponding fast ion density profiles. In addition to mode potential, HIBP measurements can also provide the density fluctuations profile associated to the Alfvén mode [melnikov2DDistributions2022], which are very relevant when comparing with the theoretical predictions.